Solve for x: (x+3)^2/3=4
Answer by stanbon(75887)

Solve for x: (x+3)^2/3=4
(x+3) = 4^(3/2)
x+3 = 2^3
x+3 = 8
x = 5
=========
Cheers,
Stan H.